OpenStudy (anonymous):

40-{50-[2+3(4+5)]} Simplify.

OpenStudy (anonymous):

19

OpenStudy (amistre64):

40-{50-[2+3(4+5)]} 40-50+[2+3(4+5)] 40-50+2+3(4+5) 40-50+2+12+15

OpenStudy (anonymous):

40-{50-[2+3(4+5)]} =40-{50-[2+3(9)]} =40-{50-[2+27]} =40-{50-[29]} =40-{21} =19

OpenStudy (amistre64):

40-50+2+12+15 40-50+4+10+15 40-50+9+10+10 40-50+9+20 60-50+9 10+9

OpenStudy (anonymous):

19